OpenStudy (anonymous):

What are the factors of 3x2 + 10x + 3? (3x − 1)(x − 3) (3x − 3)(x − 1) (3x + 1)(x + 3) (3x + 3)(x + 1)

OpenStudy (anonymous):

I would multiply the binomials together until you get the expression that's in the question but that's just me.

OpenStudy (anonymous):

First Outside Inside Last

OpenStudy (anonymous):

I did but i dont get + 10x

OpenStudy (anonymous):

10x comes from 3x * 3 + x * 1

OpenStudy (anonymous):

Outside and inside, look at the third one, there's a 3 and a 3x on the outside and a 1 and a x on the inside. Multiply the 3 and the 3x, then multiply the 1 and the x and add the two together. 9x+x=10x

OpenStudy (anonymous):

9*1=9 9+1=10 3x^2+9x+x+3 3x(x+3)+1(x+3) (3x+1)(x+3)